OpenStudy (anonymous):

Can someone help meh with a question :D A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

@BetterOffLearningThenNot

OpenStudy (anonymous):

Yo

OpenStudy (anonymous):

That'd be great. I'm just not the best at equations... Just tryin' to get this question done so i can go eat xD

OpenStudy (anonymous):

*paste

OpenStudy (anonymous):

and then click the first one

OpenStudy (anonymous):

I tried that... Didnt really help. Sorry... I'm sort of simple minded when it comes to Algebra.

OpenStudy (anonymous):

oh maybe zpupster can help

OpenStudy (anonymous):

sorry i could not help @BetterOffLearningThenNot

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Oh right. I remember... Alrighty. Just tell me what to do.. And i hope i can solve it.

OpenStudy (anonymous):

Oh its totally okay man! Hey aleast you tried thats what counts.

OpenStudy (anonymous):

I'm just confused when i try to do it. For some reason...

OpenStudy (anonymous):

I don't really know how you got to that point in the question there...

OpenStudy (anonymous):

Multiply the first equation by -2\[-6x-16y=-24\] then the bottom equation by 3\[6x+6y=9\]

OpenStudy (anonymous):

But... How did you get the -2 to devide by... This is what im stumped on.

OpenStudy (anonymous):

Whoops multiply by.

OpenStudy (anonymous):

You need a negative number to cancel out a positive number. So think of the LCM of 3 and 2. It would be 6. In order to get six you have to multiply a positive number to one equation and a negative number to the second equation so they can cancel out.

OpenStudy (anonymous):

Yeah... But why does it have to be positive? Will the equation be un-solvable if its not positive?

OpenStudy (anonymous):

Yes because if it was positive, you would not be able to cancel out the "x" variable in the equation. When you cancel out the x you will be able to find out what "y" is \[-16y +6y=-24+9\]

OpenStudy (anonymous):

OKay. So... This is getting closer to our answer huh? Yeah sorry, I'm not good at math really XD

OpenStudy (anonymous):

It's okay haha, but yea so after you get "y" all you have to do is plug it in one of the original equations and then you will have the answer for "x"

OpenStudy (anonymous):

Okay, So ima re paste this so we can answer the question... A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

What would we have to do for part A?

OpenStudy (anonymous):

I think it would be easier to start out with canceling out the "y" variable so multiply the second equation by -4 \[-8x-8y=-12\] so after the -8y and 8y cancel out you end with \[3x-8x=12-12\]

OpenStudy (anonymous):

\[-5x=0\] should be your answer and then add 5 to the other side . x=5

OpenStudy (anonymous):

This would be part a

OpenStudy (anonymous):

Then plug 5 into the x variable in both equations and solve them\[3(5)+8y=12\] \[2(5)+2y=3\]

OpenStudy (anonymous):

part b

OpenStudy (anonymous):

Okay.... Thanks so much for helping meh

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

So we got both parts here? Can you summarize it real quick for me? Just so i know wha to put for the answer?